# devoted to memory of Alexei Alexandrovich Starobinsky

### v10

@@ I Introduction
We tested the local thermodynamics and the bulk-surface correspondence in the extension of general relativity – the $f({\cal R})$ gravity. The $f({\cal R})$ gravity in terms of the Ricci scalar ${\cal R}$ is one of the simplest geometrical models, which describes the dark energy and de Sitter expansion of the Universe.Starobinsky1980 ; Starobinsky2007 ; Felice2010 ; Clifton2012 ; Nojiri2010wj ; Nojiri2017ncd ; Odintsov2023b It was used to construct an inflationary model of the early Universe – the Starobinsky inflation, which is controlled by the ${\cal R}^{2}$ contribution to the effective action. This class of models, $f({\cal R})\propto{\cal R}-{\cal R}^{2}/M^{2}$, was also reproduced in the so-called $q$-theory,KlinkhamerVolovik2008c where $q$ is the 4-form field introduced by HawkingHawking1984 for the phenomenological description of the physics of the deep (ultraviolet) vacuum (here the sign convention for ${\cal R}$ is opposite to that in Ref. Starobinsky2007 ). The Starobinsky model is in good agreement with the observations. However, despite the observational success, the theory of Starobinsky inflation is still phenomenological. Due to a rather small mass scale $M$ compared with the Planck scale it is difficult to embed the model into a UV complete theory.Chaichian2023 ; Brinkmann2023 ; Lust2023
@@ II.4 Accelerating detector: is there connection between the local process and Unruh radiation?
One may consider two independent processes related to the accelerating detector. One of them is the Unruh radiation, which could be linked to this apparent event horizon in the Rindler spacetime and to the corresponding Unruh temperature $T_{U}=a/2\pi$, where $a$ is acceleration.Unruh1976 The second one is the local process of ionization of the accelerating atom, which has no relation to the event horizon.Volovik2022e Similar to Eq.(4), the trajectory $p_{x}(x)$ of the radiated electron is obtained from the classical equation
@@
In principle, one can introduce the effective temperature, which is similar to the effective temperature $T_{\rm eff}\sim El$ of hopping electrons in a strong electric field $E$, where $l$ is the localization length.Shklovskii1992 ; Shklovskii2024 In our case the analog of the localization length is $l=1/\sqrt{m\epsilon_{0}}$. But otherwise the local process of ionization of the accelerated atom is certainly non-thermal. Since $\epsilon_{0}\ll m$, the rate of ionization in Eq.(7) essentially exceeds the ionization rate in any thermal process, which involves acceleration, $w_{\rm thermal}\sim\exp(-\gamma\epsilon_{0}/a)$. This is very different from the ionization of the atom in the de Sitter environment, which looks thermal. This demonstrates the uniqueness of the de Sitter spacetime in producing the thermal bath with local temperature.
@@ II.7 Hubble volume vs the volume of Universe, and thermal fluctuations of de Sitter state
Here $\chi_{\rm vac}$ is the vacuum compressibility – the compressibility of the fully equilibrium Minkowski vacuum with $\epsilon_{\rm vac}=-P_{\rm vac}=0$. As for the quantum fluctuations, their contribution to the vacuum energy density is typically on the order of $M_{\rm Pl}^{4}$, where $M_{\rm Pl}$ is the Planck mass. But in the equilibrium vacuum this contribution is cancelled by the trans-Planckian degrees of freedom due to thermodynamic Gibbs-Duhem relation.KlinkhamerVolovik2008 ; Volovik2013 On the other hand, the vacuum compressibility is determined by the Planck energy scale, $\chi_{\rm vac}\sim 1/M_{\rm Pl}^{4}$.
@@ III.1 Multi-metric gravity
Then the Universe can be seen as the system of $N$ sub-Universes, each with its own gravitational coupling $K_{n}$, cosmological constant $\Lambda_{n}$ and metric $g_{\mu\nu(n)}$. The action may include also $N$ fermionic speciesFrogattNielsen1991 described by $N$ independent tetrad fields $e^{a(n)}_{\mu}$:
@@
$\Theta^{a(n)}=\frac{i}{2}\left[\bar{\Psi}^{(n)}\gamma^{a}D_{\mu}\Psi^{(n)}-D_{\mu}\bar{\Psi}^{(n)}\gamma^{a}\Psi^{(n)}\right]dx^{\mu}\,,$ (15)
or by independent fünfbein fields.ObukhovVolovik2024
@@
$e^{a(n)}_{\mu}=\left<\Theta^{a(n)}\right>\,,$ (16)
while metric is the quartet of fermions.Volovik2024q
The heat exchange between the sub-Universes leads to their equilibration with formation of the common expansion rate and thus the common temperature.
@@ III.2 Heat exchange in multi-metric gravity
We consider first the system of two Universes, assuming that in each of them the entropy of horizon obeys the area law, and show that the maximum entropy corresponds to the state in which both states have the same expansion rate.
@@ VI From de Sitter thermodynamics to de Sitter decay
To describe the decay of the vacuum due to activation and thermalization of matter, the extension of the Starobinsky analysis of the vacuum decayStarobinsky1994 ; Starobinsky1996 ; Starobinsky1997 ; Starobinsky2023 is needed. Here we consider the following scenario.
The thermal exchange between the vacuum and the excited matter generates the thermal relativistic gas. The temperature of relativistic gas tends to approach the temperature $T=H/\pi$ of the de Sitter heat bath. Then the matter energy density $\epsilon_{M}$ tends to approach the value $\epsilon_{M}\sim T^{4}$, i.e. due to the heating in the de Sitter thermal bath the matter energy density tends to approach the local thermal equilibrium, $\epsilon_{M}\rightarrow bH^{4}$. Here the dimensionless parameter $b$ depends on the number of massless relativistic fields. For example, $b=7N_{F}/120\pi^{2}$ for $N_{F}$ species of massless Weyl fermions.
@@
Here we use for simplicity the conventional general relativity with $\epsilon_{\rm vac}=\Lambda=-\frac{1}{2}K_{0}{\cal R}$. This phenomenological description of the energy exchange between vacuum and matter does not depend on the details of the microscopic (UV) theory, and requires only the condition for slow variation of the Hubble parameter, $\dot{H}\ll H^{2}$.
@@
Thus the thermal character of the de Sitter state determines the process of its decay. The obtained power law decay of $H$ in Eq.(51) was also found in Refs. Padmanabhan2003 ; Padmanabhan2005 ; Markkanen2018 ; Markkanen2018a ; Roman2020 ; Gong2021 , although using different approaches. In the Padmanabhan modelPadmanabhan2003 ; Padmanabhan2005 the de Sitter horizon is considered as the photosphere with the radiative luminosity $dE/dt\propto T^{4}A_{H}$, where $A_{H}=4\pi/H^{2}$ is the area of horizon. Since the energy of the Hubble volume is $E\sim M_{\rm Pl}^{2}/H$, and $T^{4}A_{H}\sim H^{2}$, this leads to Eq.(52) for the vacuum energy density.
In Refs. Markkanen2018 ; Markkanen2018a ; Roman2020 the parameter $t_{0}$ is related to the initial value of the Hubble parameter at the beginning of inflation at $t=0$:
@@
This $H(t=0)$ corresponds to the scaleron mass $M$ in Starobinsky inflation. The time $t_{0}\sim E_{\rm Pl}^{2}/H^{3}_{t=0}$ is called the quantum breaking time of space-times with positive cosmological constant.Dvali2019 ; Dvali2022 As was mentioned by Padmanabhan,Padmanabhan2005 in his model the late time cosmological constant is independent of the initial value, see Eq.(52) at $t\gg t_{0}$.
@@ VII Conclusion
The local thermodynamics of the de Sitter state in the Einstein gravity gives rise to the Gibbons-Hawking area law for the total entropy inside the cosmological horizon. Here we extended the consideration of the local thermodynamics to the $f({\cal R})$ gravity. We obtained the same area law, but with the modified gravitational coupling $K=df/d{\cal R}$, which is in agreement with the global thermodynamics. This supports the suggestion that the de Sitter vacuum is the thermal state with the local temperature $T=H/\pi$, and that the local thermodynamics is based on the thermodynamically conjugate gravitational variables $K$ and ${\cal R}$. The variable $K$ plays the role of the chemical potential, which is constant in the thermal equilibrium.
The local temperature $T=H/\pi$ has the definite physical meaning. It is temperature, which is experienced by the external object in the de Sitter environment. In particular, this temperature determines the local activation processes, such as the process of ionization of an atom in the de Sitter environment. The nonzero local temperature of the de Sitter state suggests the thermal instability of this state due to the thermalization of matter. The process of thermalization of matter with the corresponding decay of the vacuum energy density determines the quantum breaking time of the space-times with positive cosmological constant. The thermalization of matter by the de Sitter heat bath leads to the power law decay of the vacuum energy density, which is described by two phenomenological parameters. This behaviour is rather universal, and in particular it reproduces the result of the Padmanabhan model.Padmanabhan2005
The connection between the bulk entropy of the Hubble volume, and the surface entropy of the cosmological horizon suggests a kind of the bulk-surface correspondence, which may have the holographic origin.Bekenstein1981 ; Verlinde2023 ; Milekhin2023 It would be interesting to check this correspondence using the more general extensions of the Einstein gravity and also different types of the generalized entropy.Odintsov2023 ; Odintsov2023a ; Odintsov2022 ; Odintsov2021
We also discussed the thermodynamics of de Sitter in the frame of the statistical ensemble of the multi-metric gravities. The heat exchange between different ”sub-Universes” in the ensemble leads to the common de Sitter expansion with common temperature $T=H/\pi$.
Application of the local thermodynamics to the entropy of the Schwarzschild black hole was also considered. We obtained the Bekenstein-Hawking entropy of black hole from the negative entropy of the contracting de Sitter core of the gravastar.

### v12

@@ I Introduction
We tested the local thermodynamics and the bulk-surface correspondence in the extension of general relativity – the $f({\cal R})$ gravity. The $f({\cal R})$ gravity in terms of the Ricci scalar ${\cal R}$ is one of the simplest geometrical models, which describes the dark energy and de Sitter expansion of the Universe.Starobinsky1980 ; Starobinsky2007 ; Felice2010 ; Clifton2012 ; Nojiri2010wj ; Nojiri2017ncd ; Odintsov2023b It was used to construct an inflationary model of the early Universe – the Starobinsky inflation, which is controlled by the ${\cal R}^{2}$ contribution to the effective action. This class of models, $f({\cal R})\propto{\cal R}-{\cal R}^{2}/M^{2}$, was also reproduced in the so-called $q$-theory,KlinkhamerVolovik2008d ; KlinkhamerVolovik2008c where $q$ is the 4-form field introduced by HawkingHawking1984 for the phenomenological description of the physics of the deep (ultraviolet) vacuum (here the sign convention for ${\cal R}$ is opposite to that in Ref. Starobinsky2007 ). The Starobinsky model is in good agreement with the observations. However, despite the observational success, the theory of Starobinsky inflation is still phenomenological. Due to a rather small mass scale $M$ compared with the Planck scale it is difficult to embed the model into a UV complete theory.Chaichian2023 ; Brinkmann2023 ; Lust2023
@@ II.4 Accelerating detector: is there connection between the local process and Unruh radiation?
One may consider two independent processes related to the accelerating detector. One of them is the Unruh radiation, which could be linked to this apparent event horizon in the Rindler spacetime and to the corresponding Unruh temperature $T_{U}=a/2\pi$, where $a$ is acceleration.Unruh1976 ; Unruh2019 ; Unruh2022 The second one is the local process of ionization of the accelerating atom, which has no relation to the event horizon.Volovik2022e Similar to Eq.(4), the trajectory $p_{x}(x)$ of the radiated electron is obtained from the classical equation
@@
In principle, one can introduce the effective temperature, which is similar to the effective temperature $T_{\rm eff}\sim El$ of hopping electrons in a strong electric field $E$, where $l$ is the localization length.Shklovskii1992 ; Shklovskii2024 In our case, the analog of the localization length is $l=1/\sqrt{m\epsilon_{0}}$. But otherwise the local process of ionization of the accelerated atom is certainly non-thermal. Since $\epsilon_{0}\ll m$, the rate of ionization in Eq.(7) essentially exceeds the ionization rate in any thermal process, which involves acceleration, $w_{\rm thermal}\sim\exp(-\gamma\epsilon_{0}/a)$. This is very different from the ionization of the atom in the de Sitter environment, which looks thermal. This demonstrates the uniqueness of the de Sitter spacetime in producing the thermal bath with local temperature.
@@ II.7 Hubble volume vs the volume of Universe, and thermal fluctuations of de Sitter state
Here $\chi_{\rm vac}$ is the vacuum compressibilityKlinkhamerVolovik2008 – the compressibility of the fully equilibrium Minkowski vacuum with $\epsilon_{\rm vac}=-P_{\rm vac}=0$.
Note the main difference between the thermal fluctuations and quantum fluctuations. The contribution of the quantum fluctuations of the relativistic quantum fields to the vacuum energy density is typically on the order of $M_{\rm Pl}^{4}$, where $M_{\rm Pl}$ is the Planck mass. But in the fully equilibrium vacuum state this contribution is cancelled by the ultraviolet trans-Planckian degrees of freedom due to the thermodynamic Gibbs-Duhem relation.KlinkhamerVolovik2008 ; Volovik2013 This cancellation is universal, being valid both for the relativistic vacuum states and for the non-relativistic grounds states of the condensed matter systems. The contribution of thermal fluctuations to the vacuum energy is in the range of the infrared physics, which is determined by temperature $T$.
The vacuum compressibility is not constrained by the thermodynamic laws and is determined by the ultraviolet physicsKlinkhamerVolovik2008 with its Planck energy scale, $\chi_{\rm vac}^{-1}\sim M_{\rm Pl}^{4}$. In this relation it is similar to the gravitational coupling, which is also determined by the Planck scale, $K\sim M_{\rm Pl}^{2}$. On the other hand, the temperature corrections to $\chi_{\rm vac}^{-1}$ and $K$ as well as the Casimir corrections are within the range of the infrared physics, see Ref. VolovikZelnikov2003 for the universal temperature correction to the gravitational coupling $K$. The contribution to the compressibility from infrared range was discussed in Refs.BLHu2022 ; BLHu2024 ; Widom1998 . The negative contributions to compressibility obtained in these papers do not violate the stability of the quantum vacuum, since these contributions represent the corrections, which are small compared to the main value of the vacuum compressibility, $\chi_{\rm vac}^{-1}\sim M_{\rm Pl}^{4}$.
@@ III.1 Multi-metric gravity
Then the Universe can be seen as the system of $N$ sub-Universes, each with its own gravitational coupling $K_{n}$, cosmological constant $\Lambda_{n}$ and metric $g_{\mu\nu(n)}$.
Following Frogatt and NielsenFrogattNielsen1991 one can introduce $N$ independent tetrad fields $e^{a(n)}_{\mu}$ for $N$ fermionic species. In this multi-tetrad gravity one has:
${\cal L}_{n}=e_{abcd}\left(K_{n}\,R^{ab(n)}\wedge e^{c(n)}\wedge e^{d(n)}+\Lambda_{n}\,e^{a(n)}\wedge e^{b(n)}\wedge e^{c(n)}\wedge e^{d(n)}\right)\,,$ (14)
with the corresponding action for the fermionic species:
@@
$\Theta^{a(n)}=\frac{i}{2}\left[\bar{\Psi}^{(n)}\gamma^{a}D_{\mu}\Psi^{(n)}-D_{\mu}\bar{\Psi}^{(n)}\gamma^{a}\Psi^{(n)}\right]dx^{\mu}\,.$ (16)
This can be extended to the multi-fünfbein gravity, where instead of the tetrad fields the Dirac fermions are described by the rectangular vielbein (fünfbein).ObukhovVolovik2024
@@
$e^{a(n)}_{\mu}=\left<\Theta^{a(n)}\right>\,.$ (17)
In this approach, the metric is the quartet of fermions. In principle, the so called vestigial gravity can be realized, in which the bilinear combination of fermions in Eq.(17) is zero, $e^{a(n)}_{\mu}=0$, while the metric – the quartet of fermions – is nonzero:Volovik2024q
$g_{\mu\nu(n)}=\eta_{ab}\left<\Theta_{\mu}^{a(n)}\Theta_{\nu}^{b(n)}\right>\,.$ (18)
@@ III.2 Heat exchange in multi-metric gravity
The heat exchange between the sub-Universes leads to their equilibration with formation of the common expansion rate and thus the common temperature. We consider first the system of two Universes, assuming that in each of them the entropy of horizon obeys the area law, and show that the maximum entropy corresponds to the state in which both states have the same expansion rate.
@@ VI From de Sitter thermodynamics to de Sitter decay
To describe the decay of the vacuum due to activation and thermalization of matter, the extension of the Starobinsky analysis of the vacuum decayStarobinsky1994 ; Starobinsky1996 ; Starobinsky1997 ; Starobinsky2023 is needed. Especially the revolutionary stochastic inflation approach pioneered by StarobinskyStarobinsky1982 ; Starobinsky1986 is extremely useful, although it requires some modifications.Pattison2019 ; Cruces2022 ; Cruces2022b ; Shellard2024 This also includes the so-called separate universe approach, which is somewhat similar to the multi-metric gravity discussed in Sec. III. Here we consider the simple phenomenological scenario based on the energy exchange between the thermal de Sitter and the thermal matter.
The thermal exchange between the de Sitter heat bath and the excited matter generates the thermal relativistic gas. The temperature of relativistic gas tends to approach the temperature $T=H/\pi$ of the de Sitter background. Then the energy density of this matter $\epsilon_{M}$ tends to approach the value $\epsilon_{M}\sim T^{4}$. In terms of the Hubble parameter, $\epsilon_{M}\rightarrow bH^{4}$, were the dimensionless parameter $b$ depends on the number of massless relativistic fields. For example, $b=7N_{F}/120\pi^{2}$ for $N_{F}$ species of massless Weyl fermions.
@@
Here we use for simplicity the conventional general relativity with $\epsilon_{\rm vac}=\Lambda=-\frac{1}{2}K_{0}{\cal R}$. This phenomenological description of the energy exchange between the de Sitter vacuum and matter does not depend on the details of the microscopic (UV) theory, and requires only the slow-roll condition, i.e. the slow variation of the Hubble parameter, $|\dot{H}|\ll H^{2}$, or which is the same $|\dot{T}|\ll T^{2}$.
@@
Thus the thermal character of the de Sitter state determines the process of its decay. The obtained power law decay of $H$ in Eq.(53) was also found in Refs. Padmanabhan2003 ; Padmanabhan2005 ; Markkanen2018 ; Markkanen2018a ; Roman2020 ; Gong2021 , although using different approaches. In the Padmanabhan modelPadmanabhan2003 ; Padmanabhan2005 the de Sitter horizon is considered as the photosphere with the Gibbons-Hawking temperature and with the radiative luminosity $dE/dt\propto T^{4}A_{H}$, where $A_{H}=4\pi/H^{2}$ is the area of horizon. Since the energy of the Hubble volume is $E\sim M_{\rm Pl}^{2}/H$, and $T^{4}A_{H}\sim H^{2}$, this leads to Eq.(54) for the vacuum energy density. As was mentioned by Padmanabhan,Padmanabhan2005 in his model the late time cosmological constant is independent of the initial value, see Eq.(54) at $t\gg t_{0}$.
In Refs. Markkanen2018 ; Markkanen2018a ; Roman2020 the Starobinsky stochastic inflation approach has been used. The parameter $b$ therein is proportional to the number $N$ of conformal fields and the parameter $t_{0}$ is related to the initial value of the Hubble parameter at the beginning of inflation at $t=0$:
@@
This $H(t=0)$ corresponds to the scaleron mass $M$ in Starobinsky inflation. The time $t_{0}\sim E_{\rm Pl}^{2}/H^{3}_{t=0}$ is called the quantum breaking time of space-times with positive cosmological constant.Dvali2019 ; Dvali2022
@@
VI.3 de Sitter decay and Zel’dovich stiff matter
As was mentioned by Padmanabhan,Padmanabhan2005 his model leads to a late time cosmological constant in Eq.(57), which is independent of the initial value, but its value is still far too large. Can we fix this? In Sec. V.2 we obtained that the thermodynamics of de Sitter thermal bath corresponds to the Zel’dovich stiff matter with $w=1$. Let us try the stiff matter scenario using the phenomenological approach.
For the matter with $w=1$, the dissipative Hubble friction equation is $\partial_{t}\epsilon_{\rm stiff}=-6H\epsilon_{\rm stiff}$. In our case, the stiff matter analog is thermal, with linear in $T$ entropy density, and its temperature tends to approach the heat bath temperature $T=H/\pi$. Then instead of Eqs.(51) and (52) one obtains
$\partial_{t}\epsilon_{\rm stiff}=-6H(\epsilon_{\rm stiff}-\tilde{b}H^{2})\,,$ (58)
and
$\partial_{t}\epsilon_{\rm vac}=-6\tilde{b}H^{3}\,.$ (59)
This gives the following power law decay of the vacuum energy density:
$\epsilon_{\rm vac}\sim M_{\rm Pl}^{4}\left(\frac{t_{\rm Pl}}{t+t_{0}}\right)^{2}\,,$ (60)
with the reasonable value of the vacuum energy density in the present time:
$\epsilon_{\rm vac}(t=t_{\rm present})\sim\frac{M_{\rm Pl}^{2}}{t_{\rm present}^{2}}\,.$ (61)
The same behaviour of the vacuum energy was obtained in Ref.KlinkhamerVolovik2008c (see also Ref. Volovik2013 ). This was also obtained in Ref.Klinkhamer2012 , where the Polyakov scenarioPolyakov2012 of the infrared instability of the de Sitter space was discussed.
Note that in this approach, the analog of the stiff matter comes from the gravitational degrees of freedom. In this sense it has relations to Refs. Cano2024a ; Cano2024b ; DiFilippo2024 and references therein, where the role of the gravitational degrees of freedom is discussed. Since in this approach the stiff matter is on the order of vacuum energy density, $\epsilon_{\rm stiff}\sim\epsilon_{\rm vac}$, these gravitational degrees of freedom can be responsible for dark matter.Capozziello2012
So, the phenomenological approach to the vacuum energy decay may produce different power-law decays: in Eq.(60) and in Eq.(57). But it is not excluded that these two asymptotic laws correspond to different epochs.
@@ VII Conclusion
The local thermodynamics of the de Sitter state in the Einstein gravity gives rise to the Gibbons-Hawking area law for the total entropy inside the cosmological horizon. Ee extended the consideration of the local thermodynamics to the $f({\cal R})$ gravity and obtained the same area law, but with the modified gravitational coupling $K=df/d{\cal R}$. The agreement with the traditional global thermodynamics of de Sitter supports the suggestion that the de Sitter vacuum is the thermal state with the local temperature $T=H/\pi$, and that the local thermodynamics is based on the thermodynamically conjugate gravitational variables $K$ and ${\cal R}$. The variable $K$ plays the role of the chemical potential, which is constant in the thermal equilibrium. The gravitational variables modify the thermodynamic Gibbs-Duhem relation, due to which the de Sitter thermal state becomes similar to the Zel’dovich stiff matter.
The local temperature $T=H/\pi$ has the definite physical meaning. It is temperature, which is experienced by the external object in the de Sitter environment. In particular, this temperature determines the local activation processes, such as the process of ionization of an atom in the de Sitter environment. The nonzero local temperature of the de Sitter state suggests the thermal instability of this state due to the thermalization of matter by the de Sitter heat bath. The process of thermalization of matter with the corresponding decay of the vacuum energy density determines the so-called quantum breaking time of the space-times with positive cosmological constant. The thermalization of matter by the de Sitter heat bath leads to the power law decay of the vacuum energy density, which is described by two phenomenological parameters. This behaviour is rather universal, and in particular it reproduces the result of the Padmanabhan model.Padmanabhan2005 On the other hand, the analogy with the Zel’dovich stiff matter suggests another universal behaviour, which leads to the reasonable value of the vacuum energy density in the present time.
The obtained connection between the bulk entropy of the Hubble volume, and the surface entropy of the cosmological horizon suggests a kind of the bulk-surface correspondence, which may have the holographic origin.Bekenstein1981 ; Verlinde2023 ; Milekhin2023 It would be interesting to check this correspondence using the more general extensions of the Einstein gravity and also different types of the generalized entropy.Odintsov2023 ; Odintsov2023a ; Odintsov2022 ; Odintsov2021 ; Witten2023
We also discussed the thermodynamics of de Sitter in the frame of the statistical ensemble of the multi-metric gravities. The heat exchange between different ”sub-Universes” in the ensemble leads to the common de Sitter expansion with the common temperature $T=H/\pi$. Application of the local thermodynamics to the entropy of the Schwarzschild black hole was also considered. We obtained the Bekenstein-Hawking entropy of black hole from the negative entropy of the contracting de Sitter core of the gravastar object.